OpenStudy (anonymous):

show work,4 1/2+1 3/8

OpenStudy (anonymous):

\[3\div8=0.375\]\[4.5+1.375=5.875\]\[0.875=\frac{ 7 }{ 8 }\]\[5\frac{ 7 }{ 8 }\]

OpenStudy (anonymous):

Do you have any questions?